Tuesday, April 29, 2008

Weblink and APECS are back online

I have corrected the APECS and weblink problem. If you are a student, please allow 15 minutes before trying to sign-on if the system reports that your account is locked. It will unlock after 15 minutes. Thank you, SHamilton ITS Dept.

APECS crashed - working on it

APECS had a fatal crash - I am working on it currently. SHamilton ITS Dept.

Monday, April 28, 2008

APECS and Weblink are up 7:35 PM

Apecs and Weblink are up at 7:35 PM. Sorry for the inconvenience.


APECS going down for about 25 minutes to reset system

The Apecs system needs to be restarted. The process will take about 25 minutes. ITS Department

Weblink Having Problems - ITS working on it

Weblink has errors. ITS is working to correct the problem. Thank you.

Clarification on Weblink - Faculty only this AM

As a reminder, student usage is turned off until approximately noon to allow faculty free access to input grades. ITS Dept.

Weblink up 8:10 AM

The backup job from last night was still running (preventing Weblink from starting). This was corrected and Apecs and Weblink are now available. ITS Dept.

APECS/Weblink down 04/28 7:30 a.m.

APECS and Weblink are not accessible. ITS is working on that problem.

Friday, April 25, 2008

Weblink, for Students, will be turned off at 3 PM through Monday (noon?)

Weblink access for students will be turned off today at 3 PM. This is to allow faculty full access to update grades. Sometime around noon on Monday, students will be allowed to access Weblink again.

The exact time of Weblink coming online for students will be specified by the Registrar's office.

Thank you,

ITS Department

Wednesday, April 23, 2008

Titannet server restarted

The Titannet server UDM-AD-DOMAIN was briefly disconnected and restarted at 10:55am. It was brought right back up and is fully functional.

During the down time some of our users would have been disconnected from their drives G: , H: , and other network drives.

Facilities Helpdesk short service outage 04/23 8:30 a.m.

Facilities Helpdesk service will be stopped for necessary hardware maintenance. Expected service downtime is less than 10 minutes.

Thanks for your patience and understanding.


Faculty/Student/Organization Web Server Maintenance Complete

Maintenance on the faculty/student/organization web server has been completed. All services have been restored.

Tuesday, April 22, 2008

Faculty/Student/Organization Web Server Maintenance

The faculty/student/organization web server will be off line from 7:30 a.m. to 8:00 a.m. Wednesday, April 23rd for maintenance. During this time period all hosted websites and blogs will be unavailable.

Saturday, April 19, 2008

All systems operational

To the best of our knowledge, all systems appear to be up and running.


Thursday, April 17, 2008

Faculty/Student/Organization Web Server Maintenance Complete

Maintenance on the faculty/student/organization web server has been completed. All services have been restored.

Faculty/Student/Organization Web Server Maintenance

The faculty/student/organization web server will be off line from 2:00 p.m. to 2:30 p.m. Thursday, April 17th for maintenance. During this time period all hosted websites and blogs will be unavailable.

Tuesday, April 15, 2008

Internet E-Mail

Incoming mail from 12:30 am is being queued at the moment and will be delivered just after 7:30 am.

Monday, April 14, 2008

All Systems Operational

All systems are up and functioning at this time.

Weblink is Now Operational 8:49 AM

Weblink is now working normally. Sorry for the inconvenience, ITS Department

Weblink is having problems - ITS is working on it

We are working to correct the Weblink issues. ITS Department

Friday, April 11, 2008

Long Distance @ Dental Campus

Our Long Distance Carrier "Cavalier" has restored Long Distance capabilities at the Corktown Dental Campus. Extensive testing has been done to confirm that it is indeed working as it should be. All tests were completed successfully. If you have any concerns related to this outage please feel free to contact the Helpdesk.

Tuesday, April 08, 2008

Voicemail + 5 Digit Dialing

The problems with calls not being routed into Voicemail and the inability to 5 digit dial has been resolved. ATT had been working on the issue since Friday and has recently restored all Centrex Features. We are sorry for the inconvenience. If you believe that you are still experiencing a problem related to this issue please feel free to contact the Helpdesk. Thank You.

Mailborder and Incoming E-Mail

The system is presently locked and will be restarted at 7:30 am.  Any queued mail will then be delivered.

Monday, April 07, 2008

Update on voice-mail issue

As many know, the University began to experience a problem with voice-mail system this past Friday as posted on the outage blog on http://it.udmercy.edu (which is also available directly at http://udm-its.blogspot.com).  Technically, the problem lies not with the voice-mail system but the hand-off of the call from our Centrex system to our voice-mail system.  Centrex is not properly passing the calls to the voice-mail system.  One quick work-around for those who need to get into their voice-mail box, dial the full 9,993-6440 to enter directly into the voice-mail system.


Along with this hand-off problem, we are also experiencing problems using 5 digit dialing.


AT&T was on-site Friday and returned today to continue to work on the situation.  No additional information is available at this time.



Blackboard e-mail system is up and running 04/07 9:30 a.m.

Problem causing Blackboard to fail delivering messages is fixed and normal operation is resumed.

Blackboard e-mail problem 04/07 7:30 a.m.

Blackboard e-mail system is not delivering e-mail messages to end users. ITS is working to resolve this issue.

Voice Mail not working 04/07 7:30 a.m.

Saturday, April 05, 2008

All Backlogged Mail Delivered

All backlogged mail has been delivered and we have seen a significant improvement in performance.  The solutions provider has implemented a work-around while continuing to work toward a permanent solution.

Friday, April 04, 2008

SPAM Solution Update

Significant progress has been made in processing back-logged messages.  Although some SPAM has come through, we appear to be within hours of having all unprocessed messages delivered.


Blackboard outage 04/04 at 11:35

Blackboard service was interrupted shortly today at 11:35 due to system overload. Services are restarted which caused 2 minutes out of services for Blackboard users.
ITS apologizes for this inconvenience.

Voice-Mail Randomly Failing

A problem has been identified with the voice-mail system randomly failing to allow callers to leave a message.  The problem does not appear to be with the voice-mail system but with the phone service provider.  ITS Telecomm is working with AT&T to resolve this matter as quickly as possible.

Thursday, April 03, 2008

Some Encouraging SPAM/Internet E-mail News

After our trying a number of different approaches at the instructions of SonicWall (the SPAM solutions vendor) and after sending a snapshot of the activity on our spam box, SonicWall has discovered the location of a programming bug within the code that is causing slowness in the delivery and processing of internet mail.  The discovery came around 7:00 pm EST.  Since that time they have been working on a solution and we are standing by to implement it.  It appears this will drag into Friday.



Wednesday, April 02, 2008

Update on E-Mail/Spam Situation

We regret this issue is continuing to drag on but we are working to get it resolved as quickly as possible.  The problem stems somewhat from our work to get the Emergency Notification System e-mail messages automatically coming into the University and delivered to all users.


In getting our existing solution to accommodate this, we added a “front door” system in front of our SPAM solution.  Our front door system takes in all e-mail and validates sender credentials.  Messages sent from a valid sender are then transferred over for SPAM analysis.  It is this “transfer” where our problem exists.  Messages are not being transferred at the normal or acceptable rate.  Messages come across approximately at a rate of 4,000-8,000 per hour.  We should be seeing 30,000-55,000 per hour.  We are working with our solutions provider to resolve the matter.


Please stay tuned for more info on this.

Access to Mailborder

In the ongoing work to resolve the e-mail issues we have been encountering, there will be intermittent inaccessibility to the mailborder.udmercy.edu solution.  Please bear with us.

E-mail Update

Incoming emails are currently being processed by the spam solution and being delivered.